# Exploring selection biases in Fast Radio Burst (FRB) dispersion-galaxy cross-correlations with magnetohydrodynamical simulations

### v1

@@ 3.1 Insensitivity to FRB and host galaxy properties
Because, in principle, the signal in the cross-correlation arises purely from the electrons along the FRB line of sight that are clustered around foreground galaxies, it is convenient to consider two different kinds of selection effects. The first are selection effects on observational properties of the FRB and host galaxy that arise outside of the foreground galaxy redshift shell. The second are selection effects on observational properties of the FRB that arise within the galaxy redshift shell, which we will refer to as “propagation selection effects”. In this section, we consider the former and in section 3.2 we consider the latter.
3.1.1 FRB population model
We adopt a distribution of FRBs that trace the SFR of galaxies within a specified redshift shell in our fiducial model. As discussed in Section 2.1, this introduces a window function on the DM overdensity map that mixes power on different angular scales, which can be mitigated with the OQE. We test the effect of the physical placement of FRBs on the sky in Figure 4 by comparing our measurement of $C_{\ell}^{Dg}$ with the fiducial FRB sample (criteria listed in Section 2.4) to our measurement with an identical sample of FRBs except with FRBs uniformly drawn over the sky (one FRB per pixel). We find that the measurement of $C_{\ell}^{Dg}$ is consistent when drawing FRBs uniformly over the sky and when drawing them based on SFR.
3.1.2 Optical followup selection effects
@@
As done in Section 3.1.1, we test the hypothesis that the cross correlation is robust to selection effects that bias against detecting FRBs in dim hosts directly with our simulation. Although the current magnitude limit for spectroscopic redshifts extends beyond $m_{g}\sim 21$ (Khostovan et al., 2025), we consider a conservative apparent magnitude cutoff and remove all FRBs in hosts galaxies with $m_{g}>20.7$. Our results are shown in Figure 5–as expected, the cross-correlation signal is unaffected by the magnitude selection effect on the FRB host galaxy.
3.1.3 Host DM
Measurements of the FRB DM include a component from the FRB’s host galaxy and local environment, typically of the order of tens to a few hundreds of pc cm-3 (Cordes et al., 2022). For studies that are trying to measure the electron density of the IGM or intervening halos (Connor & Ravi, 2022; Connor et al., 2024; Sharma et al., 2025), the DM contribution from the host galaxy is a nuisance term that is typically forward modeled. This is further complicated by the fact that the host DM distributions of FRBs are likely to be redshift dependent Medlock et al. (2025). One advantage of FRB-DM galaxy cross-correlations is that the host DM of the FRB is not expected to correlate with the foreground galaxies, and hence the problem of assuming a functional form for the host DM contribution is avoided.
@@ 3.1.3 Host DM
3.2 Propagation selection effects
Now we turn to selection effects on observational properties of FRBs that also arise from propagation effects, namely scattering and DM. Although the host properties of an FRB contribute to both the observed scattering and DM, we ignore those effects here since we have demonstrated that the cross correlation is robust to the properties of the FRB host in Section 3.1.
3.2.1 DM dependent selection effects
To model the DM-dependent selection effects, we use a log-normal probability function based off of the selection function in CHIME/FRB Collaboration et al. (2021b)
@@ 3.2.1 DM dependent selection effects
where DM is in units of $\mathrm{pc}\,\mathrm{cm}^{-3}$ and $a$ is a free scaling parameter. The value $a=1$ roughly corresponds to the DM-dependent selection effect found in CHIME/FRB Collaboration et al. (2021b), while $a=2$ very roughly scales it down to the FRB redshift range in this study, $z_{\text{host}}<0.4$. We compare our results when applying this cut with $a=2$ to the fiducial case in Figure 7, and find that the cross-correlation is largely unaffected.
We also test two additional DM dependent selection functions that are more aggressive: one using Equation 21 with $a=5$, and another as a stepwise function cutting all FRBs with DM$>600$ pc/cm-3. The stepwise function is motivated by the limits on realtime memory buffers that can only hold a fixed amount of data and therefore a limited dispersive sweep while FRBs are searched for in backend systems (e.g. $\sim$ 1000 pc/cm-3 for CHIME (Michilli et al., 2021)).
@@ 4 Discussion and Conclusion
Limitations of our simulation framework prevent us from matching our simulated survey parameters to realistic surveys; we can only simulate patches of the sky 6 degrees in extent and ray trace up to a maxiumum redshift of 0.4. In contrast, CHIME/FRB surveys over half the sky (CHIME/FRB Collaboration et al., 2018) with its FRBs extending to redshift 1 (Shin et al., 2023). As such, our goal is not to precisely quantify the selection effects for any given survey, but to get a first sense of what selection effects are likely to be important. This will provide guidance in planning figure surveys, which will ultimately need to perform simulations matched to their survey to interpret results. Such larger-scale simulations should be feasible in the future using our ray-tracing framework an larger-box simulations such as MilleniumTNG (Pakmor et al., 2023).
[…]
@@
Our findings suggest that cross-correlation techniques remain a promising method to probe the distribution of baryons using FRB DMs, but selection effects to dispersion must be properly accounted for to avoid biased measurements. As next-generation radio telescopes such as CHORD, DSA-3000, and BURSTT begin to collect larger and more precise samples of FRBs, our results provide a framework for understanding and mitigating selection biases that may impact FRB surveys aimed at probing the large-scale distribution of baryons.
5 Data Availability
@@ acknowledgments
A.Q.C. is partially supported by the MIT UROP program. K.W.M. holds the Adam J. Burgasser Chair in Astrophysics and received support from NSF grant 2018490.

### v2

@@ III.1 Insensitivity to FRB and host galaxy properties
Because, in principle, the signal in the cross-correlation arises purely from the electrons along the FRB line of sight that are clustered around foreground galaxies, it is convenient to consider two different kinds of selection effects. The first are selection effects on observational properties of the FRB and host galaxy that arise outside of the foreground galaxy redshift shell. The second are selection effects on observational properties of the FRB that arise within the galaxy redshift shell, which we will refer to as “propagation selection effects”. In this section, we consider the former and in section III.2 we consider the latter.
III.1.1 FRB population model
We adopt a distribution of FRBs that trace the SFR of galaxies within a specified redshift shell in our fiducial model. As discussed in Section II.1, this introduces a window function on the DM overdensity map that mixes power on different angular scales, which can be mitigated with the OQE. We test the effect of the physical placement of FRBs on the sky in Figure 4 by comparing our measurement of $C_{\ell}^{Dg}$ with the fiducial FRB sample (criteria listed in Section II.4) to our measurement with an identical sample of FRBs except with FRBs uniformly drawn over the sky (one FRB per pixel). We find that the measurement of $C_{\ell}^{Dg}$ is consistent when drawing FRBs uniformly over the sky and when drawing them based on SFR.
III.1.2 Optical followup selection effects
@@
As done in Section III.1.1, we test the hypothesis that the cross correlation is robust to selection effects that bias against detecting FRBs in dim hosts directly with our simulation. Although the current magnitude limit for spectroscopic redshifts extends beyond $m_{g}\sim 21$ (Khostovan et al., 2025), we consider a conservative apparent magnitude cutoff and remove all FRBs in hosts galaxies with $m_{g}>20.7$. Our results are shown in Figure 5–as expected, the cross-correlation signal is unaffected by the magnitude selection effect on the FRB host galaxy.
III.1.3 Host DM
Measurements of the FRB DM include a component from the FRB’s host galaxy and local environment, typically of the order of tens to a few hundreds of pc cm-3 (Cordes et al., 2022). For studies that are trying to measure the electron density of the IGM or intervening halos (Connor and Ravi, 2022; Connor et al., 2024; Sharma et al., 2025), the DM contribution from the host galaxy is a nuisance term that is typically forward modeled. This is further complicated by the fact that the host DM distributions of FRBs are likely to be redshift dependent Medlock et al. (2025). One advantage of FRB-DM galaxy cross-correlations is that the host DM of the FRB is not expected to correlate with the foreground galaxies, and hence the problem of assuming a functional form for the host DM contribution is avoided.
@@ III.1.3 Host DM
III.2 Propagation selection effects
Now we turn to selection effects on observational properties of FRBs that also arise from propagation effects, namely scattering and DM. Although the host properties of an FRB contribute to both the observed scattering and DM, we ignore those effects here since we have demonstrated that the cross correlation is robust to the properties of the FRB host in Section III.1.
III.2.1 DM-dependent selection effects
To model the DM-dependent selection effects, we use a log-normal probability function based off of the selection function in CHIME/FRB Collaboration et al. (2021a)
@@ III.2.1 DM-dependent selection effects
where DM is in units of $\mathrm{pc}\,\mathrm{cm}^{-3}$ and $a$ is a free scaling parameter. The value $a=1$ roughly corresponds to the DM-dependent selection effect found in CHIME/FRB Collaboration et al. (2021a), while $a=2$ very roughly scales it down to the FRB redshift range in this study, $z_{\text{host}}<0.4$. We compare our results when applying this cut with $a=2$ to the fiducial case in Figure 7, and find that the cross-correlation is largely unaffected.
We also test two additional DM dependent selection functions that are more aggressive: one using Equation 21 with $a=5$, and another as a stepwise function cutting all FRBs with DM>600>600 pc/cm-3. The stepwise function is motivated by the limits on realtime memory buffers that can only hold a fixed amount of data and therefore a limited dispersive sweep while FRBs are searched for in backend systems (e.g. $\sim$ 1000 pc/cm-3 for CHIME (Michilli et al., 2021)).
@@ IV Discussion and Conclusion
Limitations of our simulation framework prevent us from matching our simulated survey parameters to realistic surveys; we can only simulate patches of the sky 6 degrees in extent and ray trace up to a maximum redshift of 0.4. In contrast, CHIME/FRB surveys over half the sky (CHIME/FRB Collaboration et al., 2018) with its FRBs extending to redshift 1 (Shin et al., 2023). As such, our goal is not to precisely quantify the selection effects for any given survey, but to get a first sense of what selection effects are likely to be important. This will provide guidance in planning figure surveys, which will ultimately need to perform simulations matched to their survey to interpret results. Such larger-scale simulations should be feasible in the future using our ray-tracing framework an larger-box simulations such as MilleniumTNG (Pakmor et al., 2023).
A further limitation is the resolution at which we grid the electron field, limited to 500 kpc/$h$ pixels with the computational resources available in the JupyterLab workspace provided by IllustrisTNG. This artificially cuts off the electron power spectra at scales $k\lesssim 6\,h$ Mpc-1, and suppresses the power spectrum with the pixel window function at marginally lower $k$. However, it is not expected that extragalactic electrons are strongly clustered on these scales due to feedback processes. In our study, we consider angular scales $\ell<4000$, which corresponds to physical scales k<3.7hk<3.7\,h Mpc-1 at the galaxy redshift plane. Furthermore, our study focuses not on the magnitude of power spectrum itself, but on the impact of selection effects, and there is no reason to suspect that the latter is strongly resolution dependent.
@@
Our findings suggest that cross-correlation techniques remain a promising method to probe the distribution of baryons using FRB DMs, but selection effects to dispersion must be properly accounted for to avoid biased measurements. As next-generation radio telescopes such as CHORD, DSA-2000, and BURSTT begin to collect larger and more precise samples of FRBs, our results provide a framework for understanding and mitigating selection biases that may impact FRB surveys aimed at probing the large-scale distribution of baryons.
V Data Availability
@@ acknowledgments
A.Q.C. was partially supported by the MIT UROP program for the majority of this work, and is currently supported by the Lowell Wood Endowed Fellowship of the Fannie and John Hertz Foundation. K.W.M. holds the Adam J. Burgasser Chair in Astrophysics and received support from NSF grant 2018490.